In the laboratory synthesis of water insoluble gases, the product is collected by water displacement where the gas is bubbled into a container of water and collected in a tube inverted over the tip of delivery tube. Gases collected by this method become saturated with water vapor. The volume of as O2 gas is collected by water displacement is 80.0 ml at an atmospheric pressure of 750 torr and 30OC what is the partial pressure of O2 what is the mass of oxygen collected? The saturated vapor pressure of water at 30OC is 31.82 torr.



1
Expert's answer
2021-11-04T00:32:04-0400

P1V1= P2V2

750×80=31.82×V2

60000=31.83V2

V2= 60000/31.83

= 1.885kg